Ship reports banner with links Home IODP-USIO daily/weekly reports

< April

IODP Expedition 350

Izu-Bonin-Mariana Rear Arc

Daily Science Report for 29 May 2014

LOCATION: Underway to pilot station, Yokohama, Japan

The rig crew continued breaking down and laying out drill pipe until the ship departed from Site U1436 at 0930 h. Scientists were finalizing all the science reports. Expected time of arrival at the pilot station, Yokohama, is 0900 h, and arrival dockside is 1100 h on 30 May.


Daily Science Report for 28 May 2014

LOCATION: Site U1436 (32°23.89′N, 140°21.93′E)

At 0210 h, the last core of Expedition 350, Core U1436D-9F (59.5–62.0 mbsf; 101% recovery) was received on deck. The drill string was recovered, with the bit clearing the rig floor at 0900 h. The rig crew spent the day breaking down and laying out drill pipe while the ship was kept on station. Scientists worked on finalizing all scientific reports.


Daily Science Report for 27 May 2014

LOCATION: Site U1436 (32°23.89′N, 140°21.93′E)

Cores U1436C-9F through ‑18F (40.6–70.4 mbsf) obtained 29.76 m (100% recovery) of evolved and mafic ash, lapilli-ash, lapilli, lapillistone and ash-breccia, intercalated with mud and tuffaceous mud. The sequence recovered in Hole U1436C had less coring disturbance than those obtained in Holes U1436A and U1436B, particularly over the target interval near 50 mbsf, presumably due to using the half-length piston corer more extensively. We decided to attempt to recover one more copy of the black mafic ash layer in the shortest amount of time possible. Operations in Hole U1436C ended when the bit cleared the seafloor at 1255 h. The ship moved 20 m east of Hole U1436A and Hole U1436D was initiated at 1355 h, using a center bit to wash down to 40 mbsf. Cores U1436D-2F through ‑9F (40–62.0 mbsf) obtained 22.07 m (100% recovery) of a very similar sequence of tephra and intercalated (tuffaceous) mud as was recovered in the same interval in Holes U1436A, U1436B, and U1436C. Although several cores in Hole U1436D were plagued by a large proportion of impenetrable tephra that fell into the hole from further up in the formation, presumably due to washing down to 40 mbsf, we did eventually recover excellent samples of the targeted tephra layers. The last core of Expedition 350, Core U1436D-9F, arrived on deck at 0210 h on 28 May.


Daily Science Report for 26 May 2014

LOCATION: Site U1436 (32°23.89′N, 140°21.93′E)

After returning to Site U1436, where Hole U1436A was cored to 150 mbsf at the beginning of Expedition 350, contingency operations began with the objective to APC core two or three copies of one or two thick black ash layers recovered at ~50 mbsf in Hole U1436A. Hole U1436B was positioned 20 m north of Hole U1436A and coring was initiated at 0015 h. Cores U1436B-1H through ‑14H (0–61.8 mbsf) obtained 61.79 m (100% recovery) of evolved ash with minor evolved lapilli-ash, lapilli, and ash-breccia (14.5 m), mafic ash with minor mafic lapilli/lapillistone (7.5 m), and intercalated mud and tuffaceous mud (39.8 m). The tephra layers have an average thickness of 0.12 m. After the target depth of ~62 mbsf in Hole U1436B had been reached, the bit was raised and cleared the seafloor at 1425 h. The ship moved 20 m south of Hole U1436A and Hole U1436C was initiated at 1545 h. Cores U1436C-1H through ‑8F (0–40.6 mbsf) obtained 40.62 m (100% recovery) of a very similar sequence of tephra and intercalated (tuffaceous) mud as was recovered in the same interval in Hole U1436B.


Daily Science Report for 25 May 2014

LOCATION: Site U1436 (32°23.89′N, 140°21.93′E)

After operations were prematurely terminated by the failure of the fiber optic cable for the camera needed to reenter Hole U1437E, three scientific ideas were discussed for using the operational time remaining for our expedition. Based on operational feasibility and chances of scientific merit, the decision was to return to Site U1436 and core two or three short (~70 m) holes to recover the thick black mafic ash layers recovered in Hole U1436A at ~50 mbsf. The layers were speculated to record one or two unusually large eruptions from Aogashima volcano. One of these layers is ~2 m thick and was recovered in the top of a core, so it is not clear if this is the true thickness or if some of it may have come from further up the hole. Recovering additional, hopefully better samples, would strengthen the chances of the ash layers to be understood, and perhaps allow the air fall speculation of these deposits to be corroborated or rejected.

After recovering the drill string and seafloor positioning beacon at Hole U1437E, the ship embarked on the short transit (0930–1700 h) to Site U1436. An APC/XCB bottom hole assembly was made up and deployed to the seafloor.


Daily Science Report for 24 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Cores U1437E-77R through ‑79R (1785.0–1806.5 mbsf) recovered 8.69 m (40% recovery) of lapilli-tuff (5.7 m) and lapillistone (3.0 m). At that time (0800 h), the expected safe bit life (50 h of rotation with weight on bit) had expired and a drill string round-trip was required to replace the bit. The hole was swept with 30 barrels of high viscosity mud. The bit was raised to 1708.7 mbsf with the top drive. The top drive was disengaged for the remainder of the trip, with the bit clearing the seafloor at 1320 h and the rig floor at 1645 h.

After installing the new bit, the drill string was lowered to the seafloor (1800–2200 h). At that time the camera transmission test, carried out routinely before the camera is lowered to the seafloor to guide reentry into the hole, failed to transmit a signal, indicating that the last of three fiber cables had broken. Measurements confirmed that the fiber was damaged at ~4000 m from the camera end of the cable—an interval that was never unspooled on Expedition 350. The other two fiber strands had been broken before our expedition and a replacement fiber optic cable is scheduled for installation next week in Yokohama. We lowered the camera to the seafloor anyhow to see if decompression from the spool might bring back the connectivity, without success. This prematurely ended our expedition’s operations in Hole U1437E. Hole U1437E is currently in excellent condition for logging and/or further coring operations, preferably attempted soon before the hole deteriorates. As of 2400 h, discussions on alternative plans had just begun.


Daily Science Report for 23 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Cores U1437E-72R through ‑76R (1738.4–1785.0 mbsf) obtained 13.14 m (28% recovery) of lapilli-tuff (7.0 m), and tuff-breccia (6.1 m). Observations on ~50 clasts determined that they were highly, moderately, or sparsely pyroxene or pyroxene-feldspar phyric andesite. One 3 cm interval was interpreted as being emplaced in-situ based on a jigsaw fit contact with the surrounding hyaloclastic tuff-breccia.


Daily Science Report for 22 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Rough seas forced us to suspend coring operations from 0000–0600 h. RCB coring then resumed and Cores U1437E-68R through ‑71R (1701.0–1738.4 mbsf) obtained 22.27 m (60% recovery) of tuff-breccia (7.8 m), lapilli-tuff (6.2 m), volcanic breccia (5.3 m), lapillistone (2.7 m), and tuff (0.3 m), all polymictic. The large proportion of breccia allowed more detailed observations on 140 clasts, which were all microcrystalline to fine-grained andesite. These andesites are predominantly moderately pyroxene feldspar phyric, highly pyroxene feldspar phyric, or highly amphibole-pyroxene-feldspar phyric.


Daily Science Report for 21 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Cores U1437E-61R through ‑67R (1641.1–1701.0 mbsf) obtained 37.37 m (62% recovery) of lapilli-tuff (19.7 m), tuff (11.9 m), tuff-breccia (3.2 m), and lapillistone (2.6 m), with average bed thickness of 0.4 m. Observations of ~50 clasts determined that they are all andesite, most of them moderately to highly pyroxene feldspar phyric andesite.


Daily Science Report for 20 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

After round tripping the drill string to replace the bit, reentering Hole U1437E, and lowering the bit to 1534 mbsf on the previous day, the bit was washed to the bottom of the hole (1582.7 mbsf) by 0215 h. RCB coring resumed and Cores U1437E-55R through ‑60R (1582.7–1641.1 mbsf) obtained 26.08 m (45% recovery) of lapilli-tuff (13.2 m), lapillistone (6.8 m), tuff breccia (3.5 m), tuff (1.8 m) and volcanic breccia (0.8 m).


Daily Science Report for 19 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

The expected safe bit life (50 h of rotation with weight on bit) had expired after cutting Core U1437E-54R (1573.0–1582.7 mbsf), which arrived on deck at 0005 h. The day was spent round-tripping the drill string to replace the bit. The hole was swept with high-viscosity mud before the bit was raised with the top drive from 1583–1485 mbsf. The top drive was disengaged for the remainder of the trip and the bit cleared the seafloor at 0515 h. The drilling line was slipped and cut on the rig floor as part of the general rig maintenance before the drill string was retrieved. After the bit cleared the rig floor at 1100 h, a new bit was installed and lowered to the seafloor. The camera was deployed and Hole U1437E was reentered at 1710 h. The camera was retrieved and the bit lowered to 1489 mbsf, where resistance was encountered. The bit was pulled back to 1475 mbsf to engage the top drive and washed to 1534 mbsf by 2400 h.


Daily Science Report for 18 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Cores U1437E-48R through ‑54R (1514.6–1582.7 mbsf) obtained 12.27 m (18% recovery) of thick-bedded, matrix or clast-supported, polymictic lapilli-tuff with minor intercalations of lapillistone and tuff. All clasts (>2 cm) examined more closely are andesite.

The expected safe bit life (50 h of rotation with weight on bit) has expired again with Core U1437E-54R and most of the next day will be spent round-tripping the drill string to replace the bit.


Daily Science Report for 17 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Cores U1437E-42R through ‑47R (1456.2–1514.6 mbsf) obtained 16.2 m (28% recovery) of lapillistone (8.0 m), lapilli-tuff (6.0 m), and tuff (2.2 m), with average bed thickness of 0.48 m. The largest lapilli clasts are andesite.


Daily Science Report for 16 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Cores U1437E-34R through ‑41R (1378.4–1456.2 mbsf) obtained 47.93 m (62% recovery) of tuff breccia (1.8 m), lapillistone (3.2 m), lapilli-tuff (20.7 m), tuff (16.3 m), and mudstone (4.8 m), with average bed thickness of 0.4 m. Breccia and lapilli include rhyodacite, andesite, and basalt clasts. In addition, the first igneous body of rock was recovered at Site U1437, a 1.2 m thick interval (1388.9–1390.1 mbsf) of moderately quartz-amphibole-feldspar phyric rhyodacite intrusive sheet with pepperitic margins, indicating intrusion before the sediment was lithified. Strata are generally inclined, and faults are present with dip angles of 35°–85°.


Daily Science Report for 15 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Cores U1437E-28R through ‑33R (1320.0–1378.4 mbsf) obtained 49.54 m (85% recovery) of intercalated lapilli-tuff and lapillistone with mudstone (29.4 m), tuff with lapilli and/or mudstone (13.1 m), and mudstone (7.0 m), with average bed thickness of 0.33 m. This is a significant downhole increase in the amount of coarse-grained volcaniclastics (lapilli) and average bed thickness compared to previous days’ recovery. Strata are inclined at angles up to 60°, and small scale deformational structures are pervasive in the finer grained material, including anastomosing sub-vertical normal faults up to 1.5 m in length and with up to several cm displacement. A wiper trip was conducted from 0930–1630 h (1291–1359 mbsf) to clear the hole of drill cuttings, including two sweeps with 30 and 40 barrels of high viscosity mud, respectively.


Daily Science Report for 14 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Hole U1437E was reentered with the new RCB bit at 0010 h. The bit could not easily pass 1176 mbsf, so the bit was raised to 1164 mbsf, the top drive engaged, and the hole was washed to 1206 mbsf, where another obstruction and high torque were encountered. After raising the bit back to 1174 mbsf, where normal torque was reestablished, the bit was washed down to 1290 mbsf and no fill was detected in the bottom of the hole. The hole was swept with 40 barrels of high viscosity mud and RCB coring resumed at 0915 h. Cores U1437E-25R through ‑27R (1290.8–1320.0 mbsf) obtained 13.51 m (46% recovery) of intercalated lapilli-tuff with mudstone (9.9 m) and tuff with mudstone (3.6 m) with average thickness of 0.18 m. In this interval, strata become inclined at angles up to 30°, and small scale deformational structures are pervasive, including sub-vertical normal faults up to >30 cm length and up to 1.2 cm offset within the core face, and intricate fracture networks at cm and mm scale.


Daily Science Report for 13 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Cores U1437E-23R through ‑24R (1271.5–1290.8 mbsf) obtained 6.25 m (32% recovery) of intercalated lapilli-tuff with mudstone (4.2 m), tuff (1.7 m), sandstone (0.2 m) and mudstone (0.2 m) in ~42 discrete intervals with average thickness of 0.15 m.

The drill bit life cycle (~50 h of rotation with weight on bit) expired at this time. The bit was pulled from the bottom of the hole to 1077 mbsf with the top drive engaged. High torque, a loss of ability to circulate, and up to 40 klb overpull required to raise the drill string, indicated problematic hole conditions from 1281–1242 mbsf. After circulation was re-established and torque and pull returned to normal levels, the top drive was removed and the drill string retrieved from 1077 mbsf, clearing the seafloor at 1025 h and the rig floor at 1400 h. After a few hours of rig maintenance, a new RCB bit was installed and lowered to the seafloor. At the end of the day, the camera was being lowered to the seafloor to guide reentry into Hole U1437E.


Daily Science Report for 12 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Cores U1437E-15R through ‑22R (1194.0–1271.5 mbsf) obtained 66.84 m (86% recovery) of intercalated tuff breccia (0.5 m), lapilli-tuff and lapillistone (7.0 m), tuff (6.9 m), tuffaceous sandstone (3.5 m), and tuffaceous mudstone (48.8 m), in ~440 discrete intervals with average thickness of 0.15 m. The green to gray rocks are pervasively altered. No new age determinations were possible below ~973 mbsf (6 Ma) so far.


Daily Science Report for 11 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Cores U1437E-9R through ‑14R (1142.6–1194.0 mbsf) obtained 37.22 m (72% recovery) of intercalated tuff and tuffaceous mudstone, with a tuff/mudstone ratio of ~18 and an average tuff layer thickness of ~0.03 m. Although recovery was generally excellent (two cores with >100% recovery), we also had a barrel with zero recovery, presumably as a result of losing a perfectly cut core because it could not be broken off the formation or, upon retrieval, could not be kept in the barrel by the core catcher. A de-plugger was run to ensure the core was not obstructing the bottom of the drill string.


Daily Science Report for 10 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Cores U1437E-5R through ‑8R (1110.0–1142.6 mbsf) obtained 27.25 m (84% recovery) of intercalated tuff and tuffaceous mudstone, with a tuff/mudstone ratio of ~1.2 and an average tuff layer thickness of 0.16 m. The rotation rate (and thus the coring rate) is kept low while the 95 m long bottom-hole assembly (BHA) is still within the casing extending to 1086 mbsf. Drill collars in the BHA have a larger diameter than the regular drill pipe above and may compromise the casing and/or deviate the hole from vertical if not advanced into the formation below the (un-cemented) casing shoe with care.


Daily Science Report for 9 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

After completing installation of the 1086 m long, 10.75 inch casing in Hole U1437E on the previous day, several hours were dedicated to rig maintenance. An RCB coring bit was then made up and tripped to the seafloor, which was followed by the deployment of the camera to guide reentry of the bit into Hole U1437E at 0930 h. The bit was lowered to 1048 mbsf, where the top drive was engaged, and the bit reached the bottom of the hole (1104 mbsf) at 1345 h. The hole was swept with 20 barrels of mud and coring began, with the first core arriving on deck at 1815 h. Core U1437E-4R (1104.0–1110.0 mbsf) obtained 2.87 m (48% recovery) of evolved tuff and tuffaceous mudstone.


Daily Science Report for 8 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Tripping the 1085.6 m of 10.75 inch casing into Hole U1437E continued from 651.8 mbsf to 1000.7 mbsf, where the top drive was engaged. After further lowering the casing string to 1034.7 mbsf (0145 h), fill was encountered in the hole. The pumps were engaged and the casing was washed down to 1086.7 mbsf. At that point, the 10.75 inch casing hanger latched into the 16 inch casing hanger at the reentry cone. The latch was verified with 20 klb of pull from the rig floor. The casing was released and the stinger assembly was raised to 1057.7 mbsf. The top drive was disengaged and the trip to the surface continued, with a stop at 796.7 mbsf to retrieve the camera, clearing the seafloor at 0610 h, and arriving at the rig floor at 1100 h. The underreamer and mud motor were removed and a cementing bottom-hole assembly was made up consisting of a bull nose, two stands of drill collars, the 10.75 inch running tool, and one stand of drill collars above. The cementing assembly was deployed to seafloor and, after the camera was deployed, reentered Hole U1437E at 1700 h. The running tool was landed in the reentry cone with the bit at 55.8 mbsf. Attempts to establish circulation failed, indicating that the bottom of the casing was sealed, and the cement job was aborted. The cementing string was raised and cleared the seafloor at 1830 h, and after recovering the camera, was tripped to the surface clearing the rig floor at 2235 h. This ended the deepest casing deployment in the history of the JOIDES Resolution. The remainder of the day was spent with rig maintenance.

Scientist followed the developments on the rig floor, continued to speculate on what might be found with the imminent next phase of coring below 1104 mbsf, and gathered for a talk entitled “Weird and Wonderful World of Forams.”


Daily Science Report for 7 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Assembly of the 1085.6 m long 10.75 inch casing string was completed in the early morning with a total of 85 joints. The casing hanger was installed, the running tool engaged, and the complete assembly was hung from the moon pool doors by 0730 h. A mud motor and underreamer were made up and tested. When the underreamer arms failed to open, a backup unit was installed, which performed as expected with the arms opening at a circulation rate of 35 strokes per minute. The stinger bottom-hole assembly and the running tool were made up and run into the casing by 1945 h. The casing was tripped to the seafloor, the camera was deployed, and Hole U1437E was reentered with the 10.75 inch casing at 2115 h. When the bit had passed the 16 inch casing shoe (264.1 mbsf; 2200 h), the camera system was raised above the running tool at the top of the casing string, to be lowered together with the casing string. Lowering of the casing into Hole U1437E continued and reached 651.8 mbsf by 2400 h, leaving 433.8 m to go before the 10.75 inch casing hanger could latch into the 16 inch casing hanger at the reentry cone.


Daily Science Report for 6 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

After drilling Hole U1437E to 1104 mbsf, the target depth for 10.75 inch casing, and raising the bit up into the 264 m deep 16 inch casing, a wiper trip was conducted to ensure the fresh hole was clear. The drill string was first tripped into the hole without the top drive from 259.7 to 1016.7 mbsf. The top drive was then picked up to wash and ream the hole to 1104 mbsf. Fill was encountered at 1022.7 mbsf, and high torque at 1063.7 mbsf. After raising the bit again to 1035.7 mbsf, the hole was finally washed and reamed to total depth (1104 mbsf) at 0545 h. The hole was swept with 40 barrels of high-viscosity mud and the bit raised to 1035.7 mbsf, where another 60 barrel high-viscosity mud sweep was circulated to ensure the hole was clean. The bit was then raised to 1016.7 mbsf, where the top drive was disengaged. The drill string was tripped out of the hole, clearing the seafloor at 1115 h and the rig floor at 1500 h. Assembly of the 1089 m, 10.75 inch casing string began and 60 joints (767 m) were completed by 2400 h.

Scientists continued to analyze data and edit reports and convened for a seminar talk entitled “Sudden, Strong and Strange Submarine Eruption from South Sarigan (6S eruption) in the Mariana arc.”


Daily Science Report for 5 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Drilling without coring in Hole U1437E, using a 14.75 inch bit in preparation for 10.75 inch casing, continued until the target depth of 1104 mbsf (the total depth of RCB coring in Hole U1437D) was reached at 1745 h. The hole was cleaned with a 41 barrel mud sweep before the bit was raised to 988 mbsf and the top drive was disengaged. The bit was then further raised to 18.0 mbsf and lowered again to 260.1 mbsf, near the bottom of the 16 inch casing string. The drill string remained in that position while the rig crew was slipping and cutting the drilling line. The plan is first to conduct a wiper trip to the bottom of the hole, reaming and washing from 1060 to 1104 mbsf to ensure the hole is clear for casing installation, then to install the 10.75 inch casing to 1089 mbsf.

Scientists continued to analyze data and edit reports and convened for seminars entitled “Helium isotopic variability in Western Saudi Arabia: insights into lithospheric, metasomatic, and magmatic processes,” and “Unraveling the fate of an Ordovician rear-arc in eastern Australia.”


Daily Science Report for 4 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Hole U1437E was drilled (without coring) from 833.1 to 970.6 mbsf with a 14.75 inch bit to prepare the hole for 10.75 inch casing. Once the target depth of 1104 mbsf is reached (the total depth of RCB coring in Hole U1437D), the plan is to install 10.75 inch casing to 1089 mbsf. Scientists continued to analyze data and edit reports.


Daily Science Report for 3 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

Hole U1437E was drilled (without coring) from 551.5 to 833.1 mbsf with a 14.75 inch bit to prepare the hole for 10.75 inch casing. Once the target depth of 1104 mbsf is reached (the total depth of RCB coring in Hole U1437D), the plan is to install 10.75 inch casing to 1093 m. Scientists continued to draft, edit, and finalize reports.


Daily Science Report for 2 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

After retrieving the drill string from the aborted cementing run, the bottom-hole assembly (BHA) was equipped with a 14.75 inch tricone bit to advance Hole U1437E to 1104 mbsf for the 10.75 inch casing. The bit was tripped to the seafloor, reentered Hole U1437E at 0605 h, and was further tripped to 230.7 mbsf. The camera used to guide reentry was retrieved and the top drive was engaged to wash the hole down to 266.7 mbsf, where the cement plug was encountered. The cement plug, which extended to 271.7 mbsf, was drilled out and the hole was flushed with high viscosity mud. By 2400 h, the hole was advanced to 551 mbsf.

Scientists continued to draft, edit, and finalize reports, and convened for a seminar talk entitled “Comparing Izu forearc to reararc in the Oligocene: basement of Sites U1436/792 versus the Kyushu-Palau Ridge.”


Daily Science Report for 1 May 2014

LOCATION: Site U1437 (31°47.39′N, 139°01.58′E)

At 0030 h, the top drive was picked up to lower the drill string with the 14.75 inch tricone drill bit from the top of Hole U1437E to the bottom of the hole. At 0330 h, the bit tagged the top of the cement plug at 264.7 mbsf. This was 1 m below the 16 inch casing shoe, indicating that the cement had been displaced several meters too far down the 16 inch casing and that the casing likely had not been cemented to the formation as planned. It was decided that this required a second cementing round trip. The drill string was pulled back to 230.7 mbsf, the top drive was disengaged, and the drill string was tripped to the surface, clearing the seafloor at 0440 h and arriving at the rig floor at 1100 h. The casing running tool and logging bit were made up with the bottom-hole assembly (BHA). The drill string was lowered to the seafloor and the subsea camera was deployed to guide reentry into Hole U1437E, which was completed at 1740 h. After tripping into the hole to 55.7 mbsf using the top drive, weight was applied through the drill string to ensure the casing running tool sealed the top of the casing in preparation for cement delivery. When circulation was attempted by pumping seawater into the hole and around the bottom of the 16 inch casing, a pressure increase was observed, indicating that the bottom of the casing was actually sealed off, either by the first cementing job or by the formation sealing off around the casing since. The second cementing job was therefore aborted at 1830 h. The drill string was raised to 18.1 mbsf and the top drive was disengaged. The camera system was retrieved and the drill string was tripped to the surface (clearing the seafloor at 1915 h and the drill floor at 2350 h). Our next operation will be to drill a 14.75 inch hole for 10.75 inch casing.

Scientists continued to draft, edit, and finalize reports.




About the IODP JRSO | Expeditions | Data & core samples | Tools & laboratories | Publications | Education |
Newsroom | Meetings & port calls | Employment | Contact us | Search | Site map | People | Intranet | Home

Accesibility Statement | Plug-in download | Privacy statement | ODP

For comments or questions: webmaster@iodp.tamu.edu

Copyright IODP JRSO

Modified on Friday, 30-May-2014 16:57:07 CDT.